OpenStudy (anonymous):

What are the classifications of each system? x+5y=-2 x+5y=4 A.consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

@waterineyes @iGreen @AnswerMyQuestions @animal_lover36

OpenStudy (anonymous):

See, \(1 + 2 = \) ?

OpenStudy (anonymous):

3 @waterineyes

OpenStudy (anonymous):

Now take a look at this: I say : \(1 + 2 = 5\) \(1 + 2 = 10\) Is that ever possible?

OpenStudy (anonymous):

No @waterineyes

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

Now we take a look at your question: \(\color{green}{x+5y} = -2\) \(\color{green}{x + 5y} = 4\)

OpenStudy (anonymous):

In both the equations, the term which I have colored with Green, are same.. Right??

OpenStudy (anonymous):

yes @waterineyes

OpenStudy (anonymous):

Then is that possible that they can have two different values??

OpenStudy (anonymous):

\(x + 5y\) is just like \(1 + 2\), if \(1+2\) cannot have two different values, then \(x+5y\) can have or not?

OpenStudy (anonymous):

not @waterineyes

OpenStudy (anonymous):

If there is no values, possible, for both x and y, we say the system is ??? What we say?? Do you know?

OpenStudy (anonymous):

*are

OpenStudy (anonymous):

So, we say the system is \(\text{Inconsistent System..}\)

OpenStudy (anonymous):

Now, I can say in general that: If you have: \[ax + by = c\] \[ax + by = d \quad \quad \quad \quad (c \ne d)\] Then the system is Inconsistent System and there is no value of \(x\) and \(y\) possible for which the equations are satisfied. :)

OpenStudy (anonymous):

could you help me with a few more ? @waterineyes

OpenStudy (anonymous):

And from all the users, \[\huge \color{green}{\textsf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

I can try.. :)